audreymitchell8:

Nico is saving money for his college education. He invests some money at 7​%, and ​$2300 less than that amount at 5%. The investments produced a total of ​$221 interest in 1 yr. How much did he invest at each​ rate?

Hero:

Hi @audreymitchell8, I'm going to help you setup this problem.

Hero:

Let x represent the amount of money invested at seven percent and y represent the amount total amount of money invested Then we can setup two equations as follows: (1st) x + (x - 2300) = y (2nd) .07x + .05(x - 2300) = 221

Hero:

The 1st equation represents the total amount of money invested in both accounts The 2nd equation represents the interest gained after one year.

will:

2521